### v2

@@ 1 Introduction
Our contributions are presented as follows. In Section 2, we introduce TSM, a new threshold calibration method which matches the distribution of tile-level prediction scores to a reference distribution using optimal transport (OT), and accounts for prevalence shift with importance sampling (IS). We also provide theoretical evidence of sensitivity control in particular cases. In Section 3, we demonstrate empirically that TSM controls the sensitivity across several indications and classification tasks. Our experiments also show that, contrary to existing works, TSM controls sensitivity in extremely low data and prevalence regimes, when only 5 positive samples are available for calibration.
@@ 2 Tile-Score Distribution Matching
In this section, we present our method and explain how it can be combined with Chowder [5] to target pre-specified sensitivity levels, as illustrated in Figure 2. For clarity, we restrict to sensitivity here, but the extension of the method to reach pre-specified specificity levels is straightforward by simply considering negative samples instead of positive samples.
@@
By construction, for $1\leq i\leq N$ and for $k\in\{t,c\}$, $T_{i}^{k}$ admits a positive density function on $\mathbb{T}$. For simplicity of exposition, we assume that, in each cohort, tiles are i.i.d. conditionally to the slide label. Thus, for $1\leq i,j\leq N$, $k\in\{t,c\}$ and for $l\in\{0,1\}$, $\mathbb{P}(T_{i}^{k}|Y^{k}=l)=\mathbb{P}(T_{j}^{k}|Y^{k}=l)$. This assumption is required to prove the theoretical control of sensitivity; however, we show experimentally in section 3 that sensitivity is controlled in real-life scenarios. As the tile score function $g$ is continuous, the random variables $X_{1}=g(T_{1}^{k}),\dots,X_{N}=g(T_{N}^{k})$ are also i.i.d. conditionally to the label $Y^{k}$. Denoting $\omega^{k}=\mathbb{P}(Y^{k}=1)$ the prevalence in cohort $k$, the density function of a tile score $X$ is given for $k\in\{c,t\}$ by
@@ 4 Conclusion
In this paper, we propose a novel threshold calibration method to control sensitivity in WSI binary classification tasks. Using only a handful of calibration samples, TSM outperforms existing works in terms of sensitivity control on several indications and classification tasks. The main shortcoming of TSM is its specificity to the Chowder architecture which, although frequently used for WSI analysis, remains limiting. Thus, future work includes the extension of TSM to other MIL modeling approaches [9]. Another direction of future research is the analysis of TSM in settings where the clinical prevalence of positive samples is unknown, and we must rely on estimations based on the calibration data or from the literature.
@@ 4.0.1 Acknowledgements
The results presented here are in part based upon data generated by Medipath, Cypath and the TCGA Research Network: https://www.cancer.gov/tcga.
Authors would like to thank Alexandre Filiot and Auriane Riou for their valuable help and insights to conduct the downstream evaluations, and Lucas Fidon for his feedback.

### v3

@@ 1 Introduction
Our contributions are presented as follows. In Section 2, we introduce TSM, a new threshold calibration method which matches the distribution of tile-level prediction scores to a reference distribution using optimal transport (OT), and accounts for prevalence shift with importance sampling (IS). We also provide theoretical evidence of sensitivity control in particular cases. In Section 3, we demonstrate empirically that TSM controls the sensitivity across several indications and classification tasks [^1] Our code, model and features are available at https://github.com/owkin/tsm.git.. Our experiments also show that, contrary to existing works, TSM controls sensitivity in extremely low data and prevalence regimes, when only 5 positive samples are available for calibration.
@@ 2 Tile-Score Distribution Matching
In this section, we present our method and explain how it can be combined with Chowder [5] to target pre-specified sensitivity levels, as illustrated in Figure 2. For clarity, we restrict to sensitivity, but the extension of the method to reach pre-specified specificity levels is straightforward by simply considering negative samples instead of positive samples.
@@
By construction, for $1\leq i\leq N$ and for $k\in\{t,c\}$, $T_{i}^{k}$ admits a positive density function on $\mathbb{T}$. For simplicity of exposition, we assume that, in each cohort, tiles are i.i.d. conditionally to the slide label. Thus, for $1\leq i,j\leq N$, $k\in\{t,c\}$ and for $l\in\{0,1\}$, $\mathbb{P}(T_{i}^{k}|Y^{k}=l)=\mathbb{P}(T_{j}^{k}|Y^{k}=l)$. This assumption is required to prove the theoretical control of sensitivity, yet, it may not hold in practice, as adjacent tissue samples have correlated spatial structure. We note however that the sparsity of the tiling can be controlled, for instance by extracting only a subset of tiles, serving as an effective way to reduce the spatial correlation between the considered tiles. Besides, we show experimentally in section 3 that sensitivity is controlled in real-life scenarios.
As the tile score function $g$ is continuous, the random variables $(X_{i}:=g(T_{i}^{k}))_{1\leq i\leq N}$ are also i.i.d. conditionally to the label $Y^{k}$. Denoting $\omega^{k}=\mathbb{P}(Y^{k}=1)$ the prevalence in cohort $k$, the density function of a tile score $X$ is given for $k\in\{c,t\}$ by
@@ 4 Conclusion
In this paper, we propose a novel threshold calibration method to control sensitivity in WSI binary classification tasks. Using only a handful of calibration samples, TSM outperforms existing works in terms of sensitivity control on several indications and classification tasks. The main shortcoming of TSM is its specificity to the Chowder architecture which, although frequently used for WSI analysis, remains limiting. Thus, future work includes the extension of TSM to other MIL modeling approaches [9]. Another direction of future research is the analysis of TSM in settings where the exact prevalence of positive samples is unknown. For instance, in the context of the deployment of TSM in a new hospital, the historical prevalence of the center can be used, as well as the indication prevalence from existing epidemiological data or public health statistics.
@@ 4.0.1 Acknowledgements
The results presented here are in part based upon data generated by Medipath, Cypath and the TCGA Research Network: https://www.cancer.gov/tcga. Authors would like to thank Alexandre Filiot and Auriane Riou for their valuable help and insights to conduct the downstream evaluations, and Lucas Fidon for his feedback.
[…]
Persons affiliated with Owkin own stock-options in the company (A.P., J.K., G.R., A.O.).